OpenStudy (anonymous):

there are two sections A and B in a class, having 40 and 35 pupils respectively . 90% of A and 60% of B passed. Find the percentage of passes in the two together

OpenStudy (anonymous):

Total Pupils = 40+35 = 75 (40*0.9 + 35*0.6)/75*100 = 76% of the class passed

OpenStudy (anonymous):

In more detail... 40*0.9 = 36 people in A passed 35*0.6 = 21 people in B passed Total Students = 40+35 = 75 Total Students Passed = 36+21 = 57 57/75 (*100) = 76% of the class (A and B) passed.

OpenStudy (anonymous):

thank you:)